d. Brownie sell for $2.50 apiece ; ingredients cost $1 per brownie. Use this table to recommend the number of workers he should hire, and the output of cookies he should produce, for a wage of $13.50, $10.50, $9.00. Bear in mind that the revenue produced by each worker is the net after taking account of other variable costs.

    1. Nick buys a second oven and estimates output, see table 2
  1. Table 2. Brownie output and number of workers, two ovens
    1. Workers

      Brownie output

      1

      10

      2

      24.5

      3

      38.5

      4

      52

      5

      64.75

      6

      76.75

      7

      87.75

      8

      97.75

      9

      106.75

      10

      114.75

      Again, assume ingredients cost a dollar per brownie. Use this table to recommend the number of workers he should hire, and the output of cookies he should produce, for a wage of $21, $18, $13.50.
Answer #1
Labour Output Marginal product Value of marginal product wage1 marginal profit wage2 marginal profit Wage3 Marginal profit 3
1 10 10 10*1.5$=15$ 13.5 1.5 10.5 4.5 9 6
2 19 9 9*1.5$=13.5$ 13.5 0 10.5 3 9 4.5
3 27 8 8*1.5=12 13.5 -0.5 10.5 2.5 9 3
4 34 7 7*1.5=10.5 13.5 -3.5 10.5 0 9 .05
5 40 6 6*1.5=9 13.5 -4.5 10.5 -1.5 9 0
6 45 5 5*1.5=7.5 13.5 -6.5 10.5 -3.5 9 -2.5
7 49 4 4*1.5=6 13.5 -7.5 10.5 -4 5 9 -3
8 52 3 3*1.5=4.5 13.5 -9.5 10.5 -6.5 9 -4.5
9 54 2 2*1.5=3 13.5 -10.5 10.5 -7.5 9 -6
10 55 1 1*1.5=1.5 13.5 -12 10.5 -9 9 -7.5

Net price = $2.50-$1 = $1.50

At wage $13.5, the producer will have profit up to 1 unit of labour.

At wage $10.5, the producer will have profit up to 3 units of labour.

At wage $9, the producer will have profit up to to 4 units of labour.

The second question is done in the same manner.